Two 18 cm -long thin glass rods uniformly charged to +18nC are placed side by side, 4.0 cm apart. What are the electric field strengths E1, E2, and E3 at distances 1.0 cm, 2.0 cm, and 3.0 cm to the right of the rod on the left, along the line connecting the midpoints of the two rods?

Part A

Specify the electric field strength E1.

Express your answer to two significant figures and include the appropriate units.

Part B

Specify the electric field strength E2.

Express your answer to two significant figures and include the appropriate units.

Part C

Specify the electric field strength E3.

Express your answer to two significant figures and include the appropriate units.
Answer #1

l= 18cm

Q=+18nC

E = (1/4\pi \varepsilon 0)\left \{ Q/x\sqrt{x^2+(l/2)^2} \right \}

For E1

E(left) =17.86*104 N/C

E(right) = 5.68*104 N/C

E(total) = (17.86-5.68)*104 =12.18*104 N/C

For E2

Since it is at centre, E2 will be zero

For E3

E3 = -12.18*104 N/C